OpenStudy (anonymous):

At a candy store a bag holds up to 1 pound of candy. Each jawbreaker (x) weighs 0.25 pounds and each gumball (y) weighs 0.1 pounds. Which inequality shows the solution for all combinations of the two candies that will fit in the 1-pound bag?

OpenStudy (texaschic101):

.25x + .1y <= 1

OpenStudy (anonymous):

http://gyazo.com/df82fa0f074de1ca54ace46f422a36f8 These are the answer choices.

OpenStudy (texaschic101):

My above answer is in decimal form, however, if I put it in fraction form, then it is 1/4x + 1/10y <= 1 1/10y <= -1/4x + 1 y <= (- 1/4x + 1)/(1/10) y <= (- 1/4x + 1)(10) y <= - 10/4x + 10 y <= - 5/2x + 10 your answer is B